E is a dibasic acid H2X3 titrated against F a solution of NaOH containing 1.00g of NaOH in 250cm^3 of solution. 25cm^3 portions of the base and methyl orange were used. Average volume of acid used was 24.70cm^3. Write a balanced equation for the reaction. Hence calculate the concentration of F in mol/dm^3
1
Expert's answer
2019-03-12T15:54:00-0400

Question #86086, Chemistry / General Chemistry | for completion

E is a dibasic acid H2X3 titrated against F a solution of NaOH containing 1.00g of NaOH in 250cm^3 of solution. 25cm^3 portions of the base and methyl orange were used. Average volume of acid used was 24.70cm^3. Write a balanced equation for the reaction. Hence calculate the concentration of F in mol/dm^3 


Answer:


H2X3 + 2NaOH = Na2X3 + 2H2O

0.00125. 0.0025

Cm= 0.0025/2= 0.00125 mol/dm3 acid